Claim: The standard deviation of pulse rates of adult males is

more thanmore than

11

bpm. For a random sample of

146

adult​ males, the pulse rates have a standard deviation of

12.2

bpm. Find the value of the test statistic.

The value of the test statistic is

nothing.

​(Round to two decimal places as​ needed.)
Answer #1

Test statistic formula is given as below:

Chi-square = (n - 1)*S^2/σ^2

We are given

S = 12.2

σ = 11

n = 146

Chi-square = (146 - 1)*12.2^2/11^2 = 145* 148.84/121 = 178.362

The value of the test statistic is 178.36.